OpenStudy (princesssleelee):

Given the function f(x) = The quantity of 4x minus 2, divided by 3, which of the below expressions is correct? f−1(x) = The quantity of 2 minus 4x, divided by 3. f−1(x) = The quantity of 3x minus 2, divided by 4. f−1(x) = The quantity of negative 4x minus 2, divided by 3. f−1(x) = The quantity of 3x plus 2, divided by 4.

OpenStudy (onepieceftw):

This is the equation? \[f(x) = 4x - \frac{ 2 }{ 3 }\]

OpenStudy (princesssleelee):

Negative 4? @OnePieceFTW

OpenStudy (skullpatrol):

Lesson #1 Learn how to translate words into symbols, please :-)

OpenStudy (princesssleelee):

Sorry didnt know the wording was difficult to understand.

OpenStudy (skullpatrol):



OpenStudy (princesssleelee):

I know what each looks like in number form.

OpenStudy (skullpatrol):

Is that^ what you are trying to say?

OpenStudy (solomonzelman):

Suppose you have a function: \(\large\color{#000000 }{ \displaystyle f(x)=mx+b }\) and you want to find the inverse-function. In such case, this is how you approach the problem. `Step 0:` (( Just matter of convinience, rather than solution, this is why it's step 0)) For convinience, replace f(x) with y. After doing this, you get: \(\large\color{#000000 }{ \displaystyle y=mx+b }\) `Step 1:` Switch the x and y. \(\large\color{#000000 }{ \displaystyle x=my+b }\) `Step 2:` Solve for y. \(\large\color{#000000 }{ \displaystyle x=my+b }\) \(\large\color{#000000 }{ \displaystyle x\color{red}{-b} =my+b\color{red}{-b} }\) \(\large\color{#000000 }{ \displaystyle x-b =my }\) \(\large\color{#000000 }{ \displaystyle \color{red}{\frac{\color{black}{x-b} }{m}}=\color{red}{\frac{\color{black}{my} }{m}} }\) \(\large\color{#000000 }{ \displaystyle y=\frac{x-b}{m}}\) This can also be written as, \(\large\color{#000000 }{ \displaystyle y=\frac{1}{m}x-\frac{b}{m}}\) `Step 3.5` (just the correct notation) Since your function y is the inverse function denote it in the following way, \(\large\color{#000000 }{ \displaystyle f^{-1}(x)=\frac{1}{m}x-\frac{b}{m}}\)

OpenStudy (princesssleelee):

Yes

OpenStudy (skullpatrol):

Nice answer @SolomonZelman See @princesssleelee that is why you need to FIRST learn how to translate words into symbols.

OpenStudy (solomonzelman):

Tnx skullpatrol:)

OpenStudy (princesssleelee):

Ii understand the equation and how to solve i just continuously get a mind block when trying to place the numbers accordingly.

OpenStudy (solomonzelman):

That might be due to a lack of practice. If you practice more you will get it.

OpenStudy (skullpatrol):

The reason you get a "mind block" is because you are spending too much EFFORT on the words. If you try to replace them with symbols it will be easier to think about :-)

OpenStudy (princesssleelee):

I am sure that it is a lack of practice. However I can study the notes i have, answer other problems set in the same formation, I just cannot fathom this one.... but thanks

OpenStudy (skullpatrol):

Thanks for asking :D

OpenStudy (princesssleelee):

@Pawanyadav

OpenStudy (pawanyadav):

Do you need help in above problem?

OpenStudy (pawanyadav):

OK convert 2x plus 3 whole multiply by 3 into symbols.

OpenStudy (pawanyadav):

You are not replying any so am going.

OpenStudy (princesssleelee):

Im sorry. I was just trying to look back at notes, i still do not understand this problem. im probably over thinking it